OpenStudy (ashleyisakitty):

Emilio wishes to mix peanuts worth 2.52 per pound with almonds worth 3.80 per pound to make 480 pounds of a mixture worth 3.44 per pound. How much of each should he use?

OpenStudy (ashleyisakitty):

Ive set: P=2.52 A=3.80 M=3.44 and X = lb of P and Y=lbs of almonds. I set up the equation like this: 2.52x3.8y=3.44*480. I am not sure if i am on the right track.

OpenStudy (ashleyisakitty):

I set up the equation like this: 2.52x + 3.8y = 3.44*480. I made a typo in my original equation.

OpenStudy (is3535):

x + y = 480 2.52x + 3.80y = 480(3.44)

OpenStudy (alekos):

so we're looking at the intersection of the two lines

OpenStudy (ashleyisakitty):

Oh!! Right!! That is how we set up the equation. I am going to work it out now and check back in. Thanks.

OpenStudy (is3535):

no pro